OpenStudy (anonymous):

The exact value of cos(7pi/18) is:

OpenStudy (anonymous):

is it .99?

OpenStudy (xmoses1):

I got 1.22 with one repeating

OpenStudy (mertsj):

sin(pi/9)

OpenStudy (anonymous):

the options are A. .38 B. .99 C. \[\sqrt{\frac{ 2-\sqrt{2} }{ 4 }}\] D. \[\sqrt{\frac{ 2+\sqrt{2} }{ 4 }}\]

OpenStudy (xmoses1):

@Mertsj you can take over XP

OpenStudy (anonymous):

So whats the answer?

OpenStudy (mertsj):

Do you have a scientific calculator?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

cosine pi/9 is .93969 according to my calculator so that eliminates a and b

OpenStudy (anonymous):

are you in radians or degrees

OpenStudy (mertsj):

I used 20 degrees

OpenStudy (anonymous):

wait why are you doing pi/9 when the problem is cos(17pi/8)? Im confused

OpenStudy (mertsj):

Because I made a mistake.

OpenStudy (anonymous):

oooh lol

OpenStudy (anonymous):

is that A,B,C, or D

OpenStudy (mertsj):

70 degrees

OpenStudy (mertsj):

cos 70=.3420

OpenStudy (anonymous):

thats still not an option though. im not sure whether or not to be in radians or degrees

OpenStudy (mertsj):

C

OpenStudy (mertsj):

It makes no difference.

OpenStudy (mertsj):

cos 70 degrees = cos 7pi/18 radians

OpenStudy (anonymous):

oh ok. Thank you

OpenStudy (mertsj):

yw